(b) Find the short run supply function for each firm.

The demand curve for a product is given by Q = 800 – 0.1P where P denotes price and Q denotes quantity of the product. The industry comprises large number of firms. Each firm's cost function is 

C(q) = 2000 + 500q+ 20q2

(b) Find the short run supply function for each firm.
Answer #1

Answer

the marginal cost curve of these firms is:

Maq a (2000+ 500q + 2002) T = 500 + 409 ac() MC = 500 + 409 09

The marginal cost curve is a supply curve for the firm as the firm can supply any quantity at market price.

the supply curve of each firm is

P=500+40q